# Elastic Plateau–Rayleigh instability in soft cylinders: Surface elasticity and periodic beading

### v1

@@ 5.2 Results of the numerical simulation
We start by choosing as control parameter $\eta=\widehat{\mu}$. The bifurcation diagram, reported in fig. 12, shows the amplitude of the beading pattern $\Delta r$ versus the dimensionless shear modulus $\widehat{\mu}$, where
$\Delta r=\max_{Z\in\left[0,\,2\pi R_{0}/\left(\lambda\widehat{k}_{\text{cr}}\right)\right]}r\left(R_{0},\,Z\right)-\min_{Z\in\left[0,\,2\pi R_{0}/\left(\lambda\widehat{k}_{\text{cr}}\right)\right]}r\left(R_{0},\,Z\right).$
@@
Figure 12: (Left) Bifurcation graph showing the dimensionless beading amplitude $\Delta r/R_{0}$ versus the control parameter $\widehat{\mu}$ for $\lambda=1.4$, $\widehat{\Lambda}_{s}=40$ and $\lambda_{p}=0.8$. The cylindrical configuration becomes unstable when $\widehat{\mu}$ decreases below the critical threshold $\widehat{\mu}_{\text{cr}}$. The orange triangle denotes the theoretical stability threshold obtained with the linear stability analysis. (Right) Buckled morphology obtained for the three values of $\widehat{\mu}$, corresponding to the three points A, B and C reported in the bifurcation diagram.
In particular, we notice that the cylindrical configuration remains stable as long as $\widehat{\mu}$ is greater than the bifurcation threshold. The buckled morphology exhibits the formation of bulges spaced with long, extremely thinned regions: this phenomenon is caused by an increasingly localized beading pattern due to a progressive decrease in $\widehat{\mu}$. Structures like these are typically observed in damaged axons, where a similar morphological instability occurs when the degraded cytoskeleton is squeezed by the action of the surrounding actin cortex (Datar et al.,, 2019; Riccobelli,, 2021; Fu et al.,, 2021; Dehghany et al.,, 2024).
Similarly, we study the behaviour of the cylinder by choosing $\eta=\lambda_{p}$ as control parameter, simulating the increase of surface tension induced by the change of the medium surrounding the cylinder, as in the experiments of Mora et al., (2013). In fig. 13 we show the dimensionless beading amplitude against $\lambda_{p}$.
Figure 13: (Left) Bifurcation graph showing the dimensionless beading amplitude $\Delta r/R_{0}$ versus the control parameter $\lambda_{p}$ for $\lambda=1.4$, $\widehat{\Lambda}_{s}=40$ and $\widehat{\mu}=20.5$. The cylindrical configuration becomes unstable when $\lambda_{p}$ decreases below the critical threshold $\lambda_{p}^{cr}$. The orange triangle denotes the theoretical stability threshold obtained with the linear stability analysis. (Right) Buckled morphology obtained for the three values of $\lambda_{p}$, corresponding to the three points A, B and C reported in the bifurcation diagram.
@@
Finally, we perform the continuation analysis using $\lambda$ as the control parameter. We observe different behaviors depending on $\lambda_{p}$. As shown in fig. 14, for moderate pre-stretch, we observe a supercritical bifurcation. If the body is further axially stretched, the system returns to the unbuckled cylindrical configuration. This possibility was also explored by Taffetani and Ciarletta, 2015a . On the other hand, for smaller values of $\lambda_{p}$, a subcritical pitchfork bifurcation occurs (fig. 14b). In the nonlinear regime, if the cylinder is further stretched, the instability is not suppressed, as in the previous case. Instead, a sequence of period-halving secondary bifurcations appears. Secondary bifurcations characterized by period halving are a hallmark of a possible transition to chaos (Alligood et al.,, 1998). This might be the case for systems where the cylinder becomes extremely soft due to structural damage, as reported in the experiments of Matsuo and Tanaka, (1992) on cylindrical gels and in neurons (Datar et al.,, 2019).
@@
Figure 14: (Top) Bifurcation diagram showing the dimensionless beading amplitude $\Delta r/R_{0}$ versus the axial stretch $\lambda$ for $\widehat{\mu}=0.8$, $\widehat{\Lambda}_{s}=10$, $\lambda_{p}=0.7$ (a) and $\lambda_{p}=0.6$ (b). The orange triangles denote the marginal stability thresholds obtained with the linear stability analysis. (Bottom) Beaded morphology of the buckled cylinder for $\lambda_{p}=0.6$, corresponding to the points A, B, C, and D reported in the bifurcation diagram of panel (b).
@@ 6 Final remarks
Our findings underscore the importance of incorporating strain-dependent surface tension at the interface of soft solids. Indeed, key characteristics of the elastic Rayleigh-Plateau instability, such as the finite wavelength, cannot be captured without accounting for this elastic dependence. Future efforts will be devoted to understanding the behaviour of this system when subjected to structural damage. Hydrogel filaments have been observed to develop voids following mechanical instability (Matsuo and Tanaka,, 1992), a phenomenon reminiscent of elastic cavitation. Studying this system may shed light on the structural damage occurring in neurons during neurodegenerative diseases, where similar beaded structures are observed in axons (Pullarkat et al.,, 2006; Datar et al.,, 2019).
@@ Appendix B Coefficients of the matrix $\mathsf{M}$
In this appendix we provide the explicit expressions of the components of the $2\times 2$ matrix $\mathsf{M}$ arising from the imposition of the boundary condition eq. 26c. Since the expression for $U(r)$ depends on the value of $\lambda$ (see eqs. 33 and 35) we distinguish the two cases.

### v2

@@ 5.2 Results of the numerical simulation
We start by studying the bifurcations induced in the fully nonlinear problem by variations of the dimensionless shear modulus $\widehat{\mu}$. The bifurcation diagram, reported in fig. 12, shows the amplitude of the beading pattern $\Delta r$ versus the dimensionless shear modulus $\widehat{\mu}$, where
$\Delta r=\max_{Z\in\left[0,\,2\pi R_{0}/\left(\lambda\widehat{k}_{\text{cr}}\right)\right]}r\left(R_{0},\,Z\right)-\min_{Z\in\left[0,\,2\pi R_{0}/\left(\lambda\widehat{k}_{\text{cr}}\right)\right]}r\left(R_{0},\,Z\right).$ (41)
@@
Figure 12: (Left) Bifurcation diagram showing the dimensionless beading amplitude $\Delta r/R_{0}$ (see Eq. (41)) versus the dimensionless shear modulus $\widehat{\mu}$. Here, the axial strain $\lambda$, the dimensionless surface extensibility $\widehat{\Lambda}_{s}$, and the pre-stretch parameter $\lambda_{p}$ are set to 1.4, 40 and 0.8, respectively. The cylindrical configuration becomes unstable when $\widehat{\mu}$ decreases below the critical threshold $\widehat{\mu}_{\text{cr}}$. The orange triangle denotes the theoretical stability threshold obtained with the linear stability analysis. (Right) Buckled morphology obtained for the three values of $\widehat{\mu}$, corresponding to the three points A, B and C reported in the bifurcation diagram.
In particular, we notice that the cylindrical configuration remains stable as long as $\widehat{\mu}$ is greater than the bifurcation threshold. Nearby the marginal stability threshold, the buckled morphology reproduces the sinusoidal pattern of the linear stability analysis. In the nonlinear regime, the system exhibits the formation of bulges spaced with long, extremely thinned regions: this phenomenon is caused by an increasingly localized beading pattern due to a progressive decrease in $\widehat{\mu}$. Structures like these are typically observed in damaged axons, where a similar morphological instability occurs when the degraded cytoskeleton is squeezed by the action of the surrounding actin cortex (Datar et al.,, 2019; Riccobelli,, 2021; Fu et al.,, 2021; Dehghany et al.,, 2024).
Similarly, we study the behaviour of the cylinder subjected to progressive variations of the pre-stretch parameter $\lambda_{p}$, simulating the increase of surface tension induced by the change of the medium surrounding the cylinder, as in the experiments of Mora et al., (2013). In fig. 13 we show the dimensionless beading amplitude against $\lambda_{p}$.
Figure 13: (Left) Bifurcation diagram showing the dimensionless beading amplitude $\Delta r/R_{0}$ (see Eq. (41)) versus the pre-stretch parameter $\lambda_{p}$. Here, the axial strain $\lambda$, the dimensionless surface extensibility $\widehat{\Lambda}_{s}$ and the dimensionless shear modulus $\widehat{\mu}$ are set to 1.4, 40 and 20.5, respectively. The cylindrical configuration becomes unstable when $\lambda_{p}$ decreases below the critical threshold $\lambda_{p}^{cr}$. The orange triangle denotes the theoretical stability threshold obtained with the linear stability analysis. (Right) Buckled morphology obtained for the three values of $\lambda_{p}$, corresponding to the three points A, B and C reported in the bifurcation diagram.
@@
Finally, we perform the post-buckling simulations by continuously varying the axial strain $\lambda$. We notice different behaviors depending on $\lambda_{p}$. As shown in fig. 14, for moderate pre-stretch, we observe a supercritical bifurcation. If the body is further axially stretched, the system returns to the unbuckled cylindrical configuration. This possibility was also explored by Taffetani and Ciarletta, 2015a . On the other hand, for smaller values of $\lambda_{p}$, a subcritical pitchfork bifurcation occurs (fig. 14b). In the nonlinear regime, if the cylinder is further stretched, the instability is not suppressed, as in the previous case. Instead, a sequence of period-halving secondary bifurcations appears.
@@
Figure 14: (Top) Bifurcation diagrams showing the dimensionless beading amplitude $\Delta r/R_{0}$ (see Eq. (41)) versus the axial stretch $\lambda$. Here, the dimensionless shear modulus $\widehat{\mu}$ and the dimensionless surface extensibility $\widehat{\Lambda}_{s}$ are set to 0.8 and 10, respectively, while the pre-stretch parameter $\lambda_{p}$ is equal to 0.7 (a) and 0.6 (b). The orange triangles denote the marginal stability thresholds obtained with the linear stability analysis. (Bottom) Beaded morphology of the buckled cylinder for $\lambda_{p}=0.6$, corresponding to the points A, B, C, and D reported in the bifurcation diagram of panel (b).
@@ 6 Final remarks
The potential role of secondary bifurcations in wavelength selection has been conjectured in Mora et al., (2010) and our study confirms that they play an important role far from the bifurcation threshold, leading to a cascade of period-halving events.
Our work offers a possible explanation for the onset and development of beading in soft elastic filaments and underscores the importance of incorporating strain-dependent surface tension at the interface of soft solids. Indeed, key characteristics of the elastic Rayleigh-Plateau instability, such as the finite wavelength, cannot be captured without accounting for this elastic dependence. Future efforts will be devoted to understanding the behaviour of this system when subjected to structural damage. Hydrogel filaments have been observed to develop voids following mechanical instability (Matsuo and Tanaka,, 1992), a phenomenon reminiscent of elastic cavitation. Studying this system may shed light on the structural damage occurring in neurons during neurodegenerative diseases, where similar beaded structures are observed in axons (Pullarkat et al.,, 2006; Datar et al.,, 2019). Moreover, an experimental characterization of material surface parameters would be of interest, as has been done for linearly elastic material surfaces in Heyden et al., (2022).
@@ Appendix B Coefficients of the matrix $\mathsf{M}$
In this appendix we provide the explicit expressions of the components of the $2\times 2$ matrix $\mathsf{M}$ arising from the imposition of the boundary condition eq. 27c. Since the expression for $U(r)$ depends on the value of $\lambda$ (see eqs. 34 and 36) we distinguish the two cases.